40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 20 00:59:02 2011, kiwamu, Summary, General, Weekly report 
    Reply  Mon Jun 27 10:04:29 2011, kiwamu, Summary, General, Weekly report 
       Reply  Mon Jul 4 14:27:35 2011, kiwamu, Summary, General, Weekly report 
          Reply  Mon Jul 25 11:02:19 2011, kiwamu, Summary, General, Weekly report 
             Reply  Mon Jul 25 11:04:22 2011, kiwamu, Summary, General, Weekly report 
                Reply  Mon Jul 25 11:06:38 2011, kiwamu, Summary, General, Weekly report 
                   Reply  Mon Aug 1 22:30:59 2011, kiwamu, Summary, General, Weekly report 
Message ID: 4890     Entry time: Mon Jun 27 10:04:29 2011     In reply to: 4835     Reply to this: 4936
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Weekly report 

 Summary for the week ending June 26th.  (Number of elog entries = 53)

- SUS
  A BIO installed on 1X2.
  A peak finding script was prepared for diagonalization of the OSEM input matrices
  The suspension readout coefficients were changed to have unit of [um] and [urad] in each signal.
 
- ABSL
    LWE NPRO controller was brought by Peter King.
    The I-P cuvre and beam profile was measured. Nominal current was chosen to 1.8 [A].
    The access tube between PSL and AS table was back in place.
 
-RFPD
   The REFL55 characterization was analyzed (impedance gain = 615 Ohm, shot noise intercept current = 1.59 mA )
 
- MC
   WFS1 check, the 29MHz resonance need to be adjusted.
   The MC locking gain was increased by 6 dB to avoid an oscillation at 30 kHz.
 
- LSC
  The sensing matrices were measured in DRMI configuration and PRMI configuration
 
- Fiber experiment
   QPDY_PD was repositioned to accommodate the fiber stuff on the ETMY table.
   Succeeded in introducing the IR beam into the fiber coupler.
 
- TT characterization
    Th optic bench next to MC2 was cleaned up and leveled
 
- Vent list wiki page
   A wiki page was made for the vent detailed plan.
 
- CDS
  A foton's malfunction was found. It can run correctly only on Pianosa.
  Some Dell machines were gone to Rod Luna
 
- 40m specfic safety training for the SURFs
ELOG V3.1.3-